OpenStudy (anonymous):

Simplify each expression. Use positive exponents. PLEASE SHOW WORK EVERY DETAIL. M^3 N^-6 P^0

Nnesha (nnesha):

when you there is negative exponent you need to flip the base for example \[\huge\rm x^{-2} = \frac{ 1 }{ x^2 }\]

Nnesha (nnesha):

and when there is 0 exponent of anything answer always going to be 1 for example \[\huge\rm a^0 = 1 \] \[\huge\rm 2^0 = 1 \] \[x^0 =1\]

OpenStudy (anonymous):

I have no idea how to reply on this tablet so I hope this sends. But when it comes to algebra.... I cry.. so. As simple as u can please lol

Nnesha (nnesha):

yep i also like simple explanation that's why i just gave u the "important info with examples"if you just look at the examples i gave you should be able to do ur question

OpenStudy (anonymous):

Im trying, but im not understanding. Im sorry im dumb when it comes to this

Nnesha (nnesha):

eh nope you are not :) okay so when any number or variable have "negative exponent" you need to flip it in other words "A negative exponent means divide, because the opposite of multiplying is dividing"

OpenStudy (anonymous):

Wait. N^-6 is 1/n^6 ?????

Nnesha (nnesha):

yep that's right :-)

OpenStudy (anonymous):

And p^0 =1

Nnesha (nnesha):

now what about \[\huge\rm p^0 = ??\]

Nnesha (nnesha):

yes that's right :)

Nnesha (nnesha):

so how would you writ it as a final answer :)

OpenStudy (anonymous):

How does that simplify? And setting it up is my problem

Nnesha (nnesha):

\[\huge\rm m^3 n^{-6} p^0\] this is your one question right

OpenStudy (anonymous):

What happens with m^3

Nnesha (nnesha):

that stays same because 3 is positive

OpenStudy (anonymous):

Man am I lost

Nnesha (nnesha):

nope you r fine \[\huge\rm n^{-6} = \frac{ 1 }{ n^6 }\] and \[\huge\rm p^0 = 1\] and m^3 stayz same

Nnesha (nnesha):

 so your final answer is ?